Solve for x: x - (6/x) = 1
a) 2
b) -3
c) 3, -2
d) 2, -3
e) none of these

Answer by Nate(3500)

x - 6/x = 1
xx - 6x/x = x
x^2 - 6 = x
x^2 - x = 6
(x - 1/2)^2 = 24/4 + 1/4 = 25/4
x - 1/2 = +- 5/2
x = 1/2 +- 5/2
x = 3 and x = -2